ROTATIONAL SPECTROSCOPY AS A TOOL TO INVESTIGATE INTERACTIONS BETWEEN VIBRA-TIONAL POLYADS IN SYMMETRIC TOP MOLECULES: LOW-LYING STATES OF METHYL CYANIDE

\underline{H. S. P. MÜLLER}, I. Physikalisches Institut, Univ. Köln, 50937 Köln; and MPIfR, 53121 Bonn, Germany; B. J. DROUIN, J. C. PEARSON, LINDA R. BROWN, JPL, California Inst. of Technol., Pasadena, CA 91109, USA; I. KLEINER, A. DEHAYEM, J. ORPHAL, P. CHELIN, LISA, Univ. Paris 12 & Paris 7 & CNRS, 94010 Créteil, France; R. L. SAMS, PNNL, Richland, WA 99352, USA.

Recently, interactions between vibrational polyads were studied for propyne, H_{3}C−C≡CH; in particular those between v_{10} = 1 at 30 μm with the 15 μm dyad (v_{9} = 1, v_{10} = 2),^{a}^{b} as well as between that dyad and the 10 μm tetrad (v_{5} = 1, v_{9} = v_{10} = 1, v_{10} = 3, v_{8} = 1)^{b}. Pronounced effects were caused by Δv_{10} = ±1, ΔK = 0, Δl = ±3 Fermi-type resonances at K ≈ 12. Such resonances had not been found thus far for the isoelectronic methyl cyanide, H_{3}C−C≡N, molecule despite extensive previous spectroscopic work. As methyl cyanide is also an important interstellar molecule, in particular in hot and dense molecular cores, and as it may play a role in the atmospheres of planets or of Titan, we have recorded extensive rotational and rovibrational spectra up to ∼ 1.6 THz and ∼ 1500 cm^{−1}, respectively. The present investigations focus on the v_{8} = 0, 1, and 2 states. The ν_{8} mode in methyl cyanide corresponds to the ν_{10} mode in propyne, and it is at a rather similar energy. While the infrared data pertaining to these states help to constrain their K level structures they do not reach K ≈ 14 which are perturbed most; the l = 0 component of 2 ν_{8} may be an exception. The pure rotational data on the other hand access K levels well beyond these perturbations which can be easily recognized in the spectra. Since the ν_{9} = δ(CCH) mode in propyne is missing in methyl cyanide one would expect easier assignments and analyses. However, besides Δv_{8} = ±1, ΔK = 0, Δl = ±3 Fermi-type resonances around K of 14, additional fairly strong resonances occur at similar K values which are described by Δv_{8} = ±1, ΔK = ∓2, Δl = ±1. The latter type of resonance takes even place weakly between v = 0 and v_{8} = 1; an indication for this was seen previously.^{c} The analyses of interactions between states with v_{8} ≤ 2 have been largely completed. The results will be compared with those in propyne.

aH. S. P. Müller, P. Pracna, and V.-M. Horneman, J. Mol. Spectrosc. 216 (2002) 397−407.

bP. Pracna, H. S. P. Müller, S. Klee, and V.-M. Horneman, Mol. Phys. 102 (2004) 1555−1568.

cM. Šimečková et al., J. Mol. Spectrosc. (2004) 123−126.